# Boundary crisis and long transients of the Atlantic overturning circulation mediated by an edge state

### v1

@@ V Boundary crisis: From edge state to ghost state
So far, we have investigated the global stability of the AMOC in PlaSim-LSG at constant external forcing , with the CO2 concentration set to . However, the radiative forcing of the Earth is currently undergoing rapid change as CO2 concentrations are increasing at a rate of around 0.56% per year. Consequently, the stability landscape of the Earth system is continuously evolving as a function of (see Eq. (2)). In this nonautonomous context, attractors and edge states must be viewed in a pullback or snapshot sense as they are moving in phase space subject to the change of the control .
@@
From bifurcation theory, it is well known that there may be critical forcing levels at which the global stability landscape changes qualitatively. For example, new attractors may emerge, existing ones may disappear, or attractors may switch between periodic and non-periodic behaviour. An important case are boundary crises where an attractor is annihilated by colliding with an edge state embedded in a basin boundary. If the attractor and edge states are chaotic sets, it is common that their union after the boundary crisis forms a ghost state – a chaotic state that has a long mean lifetime yet is asymptotically unstable [65, 48]. Any trajectory initialised on the ghost state will eventually diverge from it and approach an attractor, possibly after an ultralong transient. A boundary crisis involving chaotic invariant sets may be viewed as the analogue of a saddle-node bifurcation in non-chaotic systems.
V.1 AMOC stability at increased CO2 level
@@
To explore how the stability landscape of the AMOC changes under a CO2 increase, we now consider the frozen system (i.e., fixed external forcing) at a CO2 level of . Analogously to our investigation at , we run long simulations (4000 years), initialised from the ON and OFF state obtained at , respectively. Additionally, we run the edge tracking algorithm (sec. III.2), likewise initialised from the ON and OFF states at .
@@
The OFF state at persists for the 4000 years of simulation and resembles the OFF state at in terms of AMOC strength (Fig. 12d). By contrast, the simulation initialised from the former ON state eventually collapses to the OFF state after a 2700-year long transient. For the first 1300 years, this trajectory (beige line in Fig. 12d) maintains a relatively strong AMOC, though the AMOC strength begins multi-centennial oscillations reminiscent of those in [79] but growing up to in amplitude. Then, the AMOC strength abruptly declines to less than and enters a period of large AMOC oscillations that resemble those of the edge state in period and amplitude. After six cycles, the AMOC suddenly recovers and overshoots to , thereafter steeply declines again, and eventually falls to the OFF state where the trajectory remains for the final 1200 years of simulation.
@@
Dynamically, two possible situations could explain this behaviour. The basin boundary at could have moved in the phase space such that initial conditions on the ON state at 360 ppm now lie in the basin of attraction of the OFF state. Alternatively, the ON state could have disappeared entirely at , implying a monostable regime with the OFF state being the only asymptotically stable attractor. In the following, we argue for the latter possibility.
@@ V.2 Collision of ON and edge states
Despite the fact that the ON state has lost its stability at , edge tracking between the ON and OFF states is still possible for a while. This is because the former ON state is transiently stable for a few hundred years. We can thus find pairs of initial conditions that converge to a weak and, temporarily, a strong AMOC state, respectively. Running the edge tracking algorithm at produces several large AMOC oscillations that resemble the edge state dynamics at , though the AMOC minimum is initially lower and the period of around 100 years is slightly shorter than at .
After about 750 years of edge tracking, the edge trajectory interrupts its oscillatory behaviour and follows a course that is characteristic for relaxation paths from the edge state to the former ON state. Seemingly, the edge tracking algorithm loses track of the edge state and instead approaches the ON state. However, we know that the ON state is not an attractor anymore. Rather, the former ON and edge states are now an intertwined chaotic object – a ghost state [65, 66].
To see this, let us project the dynamics onto the reduced phase space spanned by the meridional and vertical SG (see section III.4). At , the OFF state has moved slightly but qualitatively remains in the same region of the reduced phase space (Fig. 12b). The former ON state and edge state now take up a larger phase space volume and are not separated anymore. The edge trajectory oscillates in a region that overlaps with the edge state at and extends further to higher meridional and lower vertical SG values. Then, it transitions to the region of the former ON state. Conversely, the trajectory initialised from the former ON state circles around the ON state region and then transitions to the edge state region, where it undergoes the same oscillations as the edge trajectory before moving to the OFF state.
Based on the phase space view taken here, we propose that the ghost state embodies the union of two interconnected phase space regions with rotational dynamics: multi-centennial oscillations near the former ON state and centennial oscillations near the former edge state. Since both regions are not separated in phase space, trajectories can chaotically switch back and forth between both oscillatory modes until they necessarily escape the ghost state and converge to the OFF state. Indeed, we observe this behaviour in simulations initiated near the ghost state, as described below.
@@ V.3 Chaotic transients
To further explore the transient dynamics in the monostable regime, we exploit simulations produced as part of the edge tracking procedure at . Specifically, we consider the ensemble of 11 simulations used for the last iteration of edge tracking before the edge trajectory jumps to the ON state region (iteration 13). These simulations are run from nearby initial conditions interpolated between one that collapsed (member 0) and another that maintained a strong AMOC (member 10) within 500 years during the previous edge tracking iteration.
The ensemble reveals a rich, non-monotonic behaviour over a period of 2000 years (Fig. 13). Initially, all trajectories undergo a spike in AMOC strength corresponding to an excursion to the former ON state region. Thereafter, the AMOC evolution varies greatly between ensemble members. While members 0 and 9 exhibit the multi-centennial mode of variability associated with the former ON state, member 3 shows repeated episodes of edge state-like oscillations. Member 2 sustains a relatively constant, strong AMOC strength for 2000 years, whereas members 8 and 9 feature more erratic behaviour with occasional spikes of high and low AMOC strength. Only four of eleven ensemble members collapse to the OFF state within 2000 years.
This demonstrates that the transient dynamics near the ghost state are essentially unpredictable and can last for thousands of years. The long lifetime of the ghost state suggests that at our model is close to the boundary crisis where the ON and edge state merged, which occurs somewhere between 360 and . Further away from the critical CO2 value, the ghost state is expected to have a shorter lifetime (see Ref. [48]). Indeed, attempting to perform edge tracking at 500 and proved unsuccessful because the model quickly diverged from the ON state, as a result of the enhanced instability of the system.
VI Role of the edge state under non-autonomous climate forcing
Our study has been focusing on the stability landscape of PlaSim-LSG at fixed external forcing, i.e., we investigated the model as an autonomous dynamical system at two different CO2 forcing levels. What can our results tell us about the transition behaviour of the AMOC in a non-autonomous context?
VI.1 Phase space trajectories under future SSP scenarios
@@
We can now inspect these simulations in the reduced phase space projection to see how their trajectories in phase space relate to the model’s stability landscape, particularly the edge state and ghost state (Fig. 14). For SSP1-2.6, the ensemble members remain in the region of the ON state. In the SSP4-6.0 scenario, the trajectories pass straight through the ghost state region, as if the oscillatory regime of the ghost state would be “invisible” to them. For SSP2-4.5, over the 1000-year simulation period, one ensemble member remains to the left of the edge state region, maintaining a strong AMOC; some trajectories travel through the lower part of the edge state region (where the AMOC is weakest, see Fig. 11a) and collapse to the OFF state; yet other ensemble members perform one or more cycles of an oscillatory motion before converging to the OFF state. These oscillations occur in the region of the edge and ghost states or to the right of it. Since the ghost state at extends to higher meridional SG values than the edge state at , it is likely that the ghost state at CO2 levels above expands to even higher meridional SG values.
These results indicate that the edge state and, beyond the boundary crisis, the corresponding ghost state play a key role in the ensemble splitting with respect to the AMOC strength, observed under the intermediate CO2 forcing. Under the low emissions scenario, the trajectories do not travel to the edge state region, while under the high emissions scenario the forcing rate is so high that the dynamical structure of the frozen system is masked. Indeed, the SSP1-2.6 scenario remains below (besides a short overshoot, see Fig. 2a), such that the ON state continues to exist. By contrast, the SSP2-4.5 and SSP4-6.0 scenarios stay above after the year 2050, such that we assume the model is in the monostable regime from then onwards.
@@ VI.2 Stochastic bifurcation in the GISS model
The diverging AMOC behaviour of ensemble members observed in PlaSim-LSG under SSP2-4.5 is intriguingly reminiscent of the so-called “stochastic bifurcation” found under the same forcing scenario in the aforementioned GISS model [52], a CMIP6 model. Can the global stability perspective presented here help explain the dynamics in GISS?
@@
We select three ensemble members of the GISS simulations under SSP2-4.5 (members 1, 7, and 10). The simulations extend until the year 2500 and show divergent AMOC behaviour (Fig. 15a). Following an initial AMOC weakening in all members, the first member starts to recover around the year 2100, the second recovers after 2200, whereas the third remains in a weak AMOC state until 2500 (but eventually recovers, see Ref. [52]). Similar to PlaSim-LSG, the weak AMOC state is characterised by a collapse of the overturning cell north of 45∘N while a weak overturning circulation is maintained south of 45∘N (Fig. 15h). In the year 1995, when CO2 levels are at and the AMOC is in the ON state, the AMOC is around 30% stronger in GISS than in PlaSim-LSG, though the meridional streamfunctions have a qualitatively similar shape (Fig. 15e, g)).
@@
To relate these simulations to our results, we likewise select three ensemble members from the SSP2-4.5 simulations with PlaSim-LSG, based on their qualitative similarity with each of the GISS ensemble member (Fig. 15c). The first member maintains a strong AMOC, the second undergoes a weakening to about followed by a recovery, and the third collapses to the OFF state. Note that the evolution of the trajectories is delayed in PlaSim-LSG compared to the GISS model, and the initial AMOC weakening is less pronounced.
@@
We now compare the reduced phase space trajectories of these simulations between the two models. Using the same definitions as for PlaSim-LSG, we compute the meridional and vertical SG in GISS based on the Atlantic zonally averaged salinity field. Due to the complexity of the GISS model, its AMOC stability landscape with respect to CO2 and the properties of potential edge states or ghost states are not known. However, we can study how the trajectories relate to the edge state and ghost state found in PlaSim-LSG.
@@
Remarkably, the reduced phase space dynamics are qualitatively similar between the two models (Fig. 15b, d). The GISS trajectories start off from a significantly lower meridional SG than the PlaSim-LSG trajectories, in line with the fact that the AMOC in GISS is significantly stronger at that time. The vertical SG values are in good agreement between the models. As CO2 forcing increases, all trajectories initially move towards larger meridional and slightly larger vertical SG values. The AMOC recovery in GISS is characterised by a counter-clockwise rotation, where the loop performs a larger excursion for the trajectory that recovers from lower AMOC values. Interestingly, the reversal of the GISS trajectory with a late recovery occurs directly in the phase space region where the edge state is located in PlaSim-LSG, following a path that resembles that of the recovering PlaSim-LSG trajectory. The collapsing trajectory in GISS skims the bottom end of the ghost state region before traveling to high meridional and low vertical SG values. This path is qualitatively similar to the collapsing PlaSim-LSG trajectory. Although the collapsing GISS simulation does not display any AMOC oscillations seen in the collapsing PlaSim-LSG simulation, there are still upward spikes in the phase space trajectory that might hint at similar, yet dampened dynamics.
@@
To summarise, we find that the splitting of the GISS ensemble occurs in the same region of the projected phase space in which the edge state is located in PlaSim-LSG. This supports the proposition that the “stochastic bifurcation” could indeed be a signature of a chaotic AMOC edge state near a boundary crisis.
@@ VII Discussion and conclusion
Traditionally, studies of climate tipping points often focus on the local dynamics near stable equilibria. Particularly, statistical EWS based on critical slowing-down measure changes in the local stability of an attractor under a quasi-adiabatic parameter drift: as the system approaches a bifurcation, the quasipotential flattens around the attractor, implying a reduction in restoring forces. Another way to look at this is that the energetic barrier imposed by the edge state diminishes towards the bifurcation, and that the quasipotential flattens around the edge state (with an opposite sign of the curvature). Closer to the crisis, the edge state becomes “stickier” [100, 54] in the sense that trajectories tend to spend longer times in its vicinity, suggesting an alternative, non-local angle on critical slowing down 444Ying-Cheng Lai, personal communication.. In fact, it seems clear that the current rate of anthropogenic greenhouse gas emissions is forcing the climate system out of a steady-state. Our results indicate that the edge state dynamics can become relevant under plausible future emissions scenarios: the boundary crisis in PlaSim-LSG occurs at CO2 levels that could be reached within two decades. Nonetheless, whether this boundary crisis is a feature of the real climate system remains unknown.
@@
This study was conducted with a climate model of intermediate complexity that inevitably relies on simplifications and neglects numerous processes of potential relevance. Therefore, the results of our investigation may be highly model-dependent and not representative of reality. Even though the AMOC edge state found in PlaSim-LSG is a physically sensible steady-state, its nonlinear dynamics might be exaggerated effects of the highly simplified parameterisations of, e.g., sea ice, oceanic convection and ocean eddies. On the other hand, we believe that PlaSim-LSG is to date the most comprehensive earth system model in which an edge state has been explicitly computed. Our results thus add a significant step towards realism to recent studies investigating edge states of the AMOC in a conceptual climate model [48] and a global ocean circulation model [57]. Furthermore, the similarity in the dynamics between PlaSim-LSG and the more complex GISS model suggests that the global stability view established here could provide key insights into the behaviour of state-of-the-art earth system models. As is increasingly clear, AMOC metastability and tipping behaviour is not restricted to simple climate models but occurs across the model hierarchy [52, 33, 78, 102]. The phase space of large models is just more challenging to explore.
@@
The key limitation for applying the edge tracking algorithm in even higher-dimensional systems is the computational cost of running long simulations. Producing 1400 years of edge trajectory in PlaSim-LSG at required around years of simulation (roughly 3000 CPU hours), giving an efficiency of 2%. However, this number depends strongly on the system, particularly how quickly trajectories diverge from the basin boundary. Improving the efficiency of edge tracking could be a valuable objective of future work.
@@
The edge tracking algorithm converges to an edge state despite the complex geometry of the basin boundary, which is typically fractal [63, 10, 48]. In our study, we observed that the basin boundary is folded along transects of interpolated initial conditions, supporting the assumption of fractality. Due to the long lifetime of the ghost state, edge tracking appears to work for multiple iterations even beyond the boundary crisis (in the monostable regime). This permits probing ghost states while also demonstrating the difficulty of precisely determining critical forcing thresholds [48]. From a modelling perspective, it is not obvious that the interpolation between initial conditions in all dynamical variables yields new initial conditions that generate numerically stable and physical trajectories. We argue that convexity of the equations governing the climate dynamics ensures that the trajectory quickly relaxes to a physical state.
@@
Using an ocean-only model, Ref. [57] found that the AMOC edge state features a less “spicy” (i.e., colder and fresher) deep North Atlantic than the attractors, as well as a higher dynamic enthalpy. Our results show the same properties for a coupled climate model. However, the atmosphere-ice-ocean interactions represented here reveal a much richer dynamics in which the upper ocean plays a more active role. Whereas Ref. [57] concluded that the most relevant regions for anticipating AMOC transitions are located in the deep sea, our results suggest that many eccentric observables are found also in the surface ocean. This could potentially be exploited for improved early warning systems of AMOC changes [99, 103].
@@
AMOC oscillations have received wide interest due to their occurrence in various climate models and potential for explaining past abrupt climate change. In the context of Dansgaard-Oeschger events [3], previous work has determined “sweet spots” for oscillations in parameter space [104]. We demonstrate a sweet spot in phase space: while the ON and OFF states do not exhibit oscillations at 360 ppm, oscillations occur near the edge state. The drivers of these unstable oscillations involve similar processes previously identified in stable oscillation mechanisms in other models [105, 106, 107]. Dynamically, the presence of unstable oscillations near the crisis might hint at the existence of a subcritical Hopf bifurcation with respect to CO2. Oscillations in the AMOC strength and other popular observables such as the freshwater export further suggest that such quantities may be poor indicators of AMOC stability in out-of-equilibrium conditions.
@@
The limited predictability of the AMOC near an instability has already been suggested by Ref. [49]. Refs. [50, 52] have recently reiterated this idea by demonstrating an ensemble splitting caused by internal variability under identical time-dependent forcing. In the context of the Snowball Earth transition, [51] found an ensemble splitting in PlaSim (without the dynamic LSG ocean). Our findings allow to understand this behaviour in terms of an edge state and, beyond the boundary crisis, a ghost state. We can thus directly link the dynamics of earth system models to fundamental concepts of dynamical systems theory that are often only explored in low-dimensional systems.

### v2

@@ V Boundary crisis: From edge state to ghost state
So far, we have investigated the global stability of the AMOC in PlaSim-LSG at constant external forcing , with the CO2 concentration set to 360 ppm. However, the radiative forcing of the Earth is currently undergoing rapid change as CO2 concentrations are increasing at a rate of around 0.56% per year. Consequently, the stability landscape of the Earth system is continuously evolving as a function of (see Eq. (2)). In this nonautonomous context, attractors and edge states must be viewed in a pullback or snapshot sense as they are moving in state space subject to the change of the control [102, 44].
@@
From bifurcation theory, it is well known that there may be critical forcing levels at which the global stability landscape changes qualitatively. For example, new attractors may emerge, existing ones may disappear, or attractors may switch between periodic and non-periodic behaviour [55]. An important case are boundary crises where an attractor is annihilated by colliding with an edge state embedded in a basin boundary [103, 104]. It has been proposed that their union after the boundary crisis forms a ghost state – a state reminiscent of the dynamics on the former attractor and edge state that has a long mean lifetime yet is asymptotically unstable [66, 48, 67]. Any trajectory initialised on the ghost state will eventually diverge from it and approach an attractor, possibly after an ultralong transient. A boundary crisis involving chaotic invariant sets may be viewed as the analogue of a saddle-node bifurcation in non-chaotic systems.
V.1 AMOC stability landscape as a function of CO2 level
@@
To explore how the stability landscape of the AMOC changes as a function of CO2, we now consider the frozen system (i.e., fixed external forcing) at two additional CO2 levels: 285 ppm (preindustrial conditions) and 460 ppm (Fig. 11). Analogously to our investigation at 360 ppm, we run long simulations (4000 years), initialised from the ON and OFF state obtained at 360 ppm, respectively. Additionally, we run the edge tracking algorithm (sec. III.2).
In the case of 285 ppm, the stability landscape qualitatively resembles the situation at : the ON and OFF states, largely unchanged in AMOC strength, are clearly separated from an oscillatory edge state (Fig. 11a,d). The edge state oscillations have a similar period but a slightly smaller amplitude compared to . Interestingly, we observed that edge tracking (initialised from the ON and OFF states at 285 ppm) is considerably more time-consuming at lower CO2, since trajectories tend to diverge more quickly from the basin boundary (40 iterations yielded about 400 years of edge trajectory instead of 1400 years obtained for ). This suggests that the edge state is more repelling at 285 ppm.
@@
At a higher CO2 level of 460 ppm, the OFF state persists for the 4000 years of simulation and resembles the OFF state at lower CO2 in terms of AMOC strength (Fig. 11f). By contrast, the simulation initialised from the former ON state eventually collapses to the OFF state after a 2700-year long transient. During the first 1300 years, this trajectory (beige line in Fig. 11f) maintains a relatively strong AMOC with multi-centennial oscillations reminiscent of those in Ref. [80], growing up to 10 Sv in amplitude. Then, the AMOC abruptly declines to less than 5 Sv and enters a period of large oscillations that resemble those of the edge state in period and amplitude. After six cycles, the AMOC suddenly recovers and overshoots to 22 Sv, thereafter steeply declines again, and eventually collapses to the OFF state where the trajectory remains for the final 1200 years of simulation.
@@
Dynamically, two possible situations could explain this behaviour. The basin boundary at 460 ppm could have moved in the state space such that initial conditions on the ON state at now lie in the basin of attraction of the OFF state. Alternatively, the ON state could have disappeared entirely at 460 ppm, implying a monostable regime with the OFF state being the only asymptotically stable attractor. In the following, we argue for the latter possibility.
@@ V.2 Collision of ON and edge states
Despite the fact that the ON state has lost its stability at 460 ppm, edge tracking between the ON and OFF states is still possible for a while. This is because the former ON state is transiently stable for a few hundred years. We can thus find pairs of initial conditions that converge to a weak and, temporarily, a strong AMOC state, respectively. Running the edge tracking algorithm at 460 ppm (initialised from the ON and OFF states at ) produces several large AMOC oscillations that resemble the edge state dynamics at 360 ppm, though the AMOC minimum is initially lower and the period of around 100 years is slightly shorter than at 360 ppm.
After about 750 years of edge tracking, the edge trajectory interrupts its oscillatory behaviour and follows a course that is characteristic for relaxation paths from the edge state to the former ON state. Seemingly, the edge tracking algorithm loses track of the edge state and instead approaches the ON state. However, we have seen that this “ON state” is not an attractor anymore. Rather, the former ON and edge states are now an intertwined chaotic object – a ghost state [66, 67].
To see this, let us project the dynamics onto the reduced state space spanned by the meridional and vertical SG (see section III.4). As CO2 increases, all states shift slightly towards larger meridional SG values (Fig. 11). The OFF and edge state display higher variability, taking up an increasing volume in the reduced state space. Strikingly, at 460 ppm, the former ON state and edge state now extend so much that they “touch” and are not separated anymore. Ocillations of the edge trajectory extend further to higher meridional and lower vertical SG values compared to the edge state at . Eventually, the edge trajectory transitions to the region of the former ON state, and the trajectory initialised from the former ON state circles around the ON state region and then transitions to the edge state region, where it undergoes the same oscillations as the edge trajectory before moving to the OFF state.
Based on the state space view taken here, we propose that the ghost state embodies the union of two interconnected state space regions with rotational dynamics: multi-centennial oscillations near the former ON state and centennial oscillations near the former edge state. Since both regions are not separated in state space, trajectories can chaotically switch back and forth between both oscillatory modes until they necessarily escape the ghost state and converge to the OFF state.
@@ V.3 Chaotic transients
To further explore the transient dynamics in the monostable regime, we exploit simulations produced as part of the edge tracking procedure at 460 ppm. Specifically, we consider the ensemble of 11 simulations used for the last iteration of edge tracking before the edge trajectory jumps to the ON state region (iteration 13). These simulations are run from nearby initial conditions interpolated between one that collapsed (member 0) and another that maintained a strong AMOC (member 10) within 500 years during the previous edge tracking iteration.
The ensemble reveals a rich transient behavior (Fig. 12). Initially, all trajectories undergo a spike in AMOC strength corresponding to an excursion to the former ON state region. Thereafter, the AMOC evolution varies greatly between ensemble members. While some members collapse after around 400-500 years, others take over 6000 years before collapsing. In fact, two members do not collapse within 7000 years of simulation. Nonetheless, we expect them to eventually collapse if the simulation would be extended. As seen from members 2 and 9, for example, the collapse can happen relatively abruptly without apparent pre-warning. During the transients, ensemble trajectories exhibit the different patterns of variability associated with the ghost state: slower, less regular oscillations of a stronger AMOC associated with the former ON state, and episodes of more rapid edge state-like oscillations.
This demonstrates that the transient dynamics near the ghost state are essentially unpredictable and can last for thousands of years. The long lifetime of the ghost state suggests that at 460 ppm our model is close to the boundary crisis where the ON and edge state merged, which occurs somewhere between 360 and 460 ppm. Further away from the critical CO2 value, the ghost state is expected to have a shorter lifetime (see Ref. [48]). Indeed, attempting to perform edge tracking at 500 and 540 ppm proved unsuccessful because the model quickly diverged from the ON state, as a result of the enhanced instability of the system.
VI Role of the edge state under nonautonomous climate forcing
Our study has been focusing on snapshots of the stability landscape of PlaSim-LSG at fixed external forcing: we investigated the model as an autonomous dynamical system at different CO2 concentrations. What can our results tell us about the transition behaviour of the AMOC in a nonautonomous context?
VI.1 State space trajectories under future SSP scenarios
@@
We can now inspect these simulations in the reduced state space projection to see how their trajectories in state space relate to the model’s stability landscape, particularly the edge state and ghost state (Fig. 13). For SSP1-2.6, the ensemble members remain in the region of the ON state. In the SSP4-6.0 scenario, the trajectories pass straight through the ghost state region, as if the oscillatory regime of the ghost state would be “invisible” to them. For SSP2-4.5, over the 1000-year simulation period, one ensemble member remains to the left of the edge state region, maintaining a strong AMOC; some trajectories travel through the lower part of the edge state region (where the AMOC is weakest, see Fig. 10a) and collapse to the OFF state; yet other ensemble members perform one or more cycles of an oscillatory motion before converging to the OFF state. These oscillations occur in the region of the edge and ghost states or to the right of it. Since the states tend to move to higher meridional SG values with increasing CO2 (Fig. 11), it is likely that the ghost state likewise expands to higher meridional SG values above 460 ppm.
These results indicate that the edge state and, beyond the boundary crisis, the corresponding ghost state play a key role in the ensemble splitting with respect to the AMOC strength, observed under the intermediate CO2 forcing. Under the low emissions scenario, the trajectories do not travel to the edge state region, while under the high emissions scenario the forcing rate is so high that the dynamical structure of the frozen system is masked. Indeed, the SSP1-2.6 scenario remains below 460 ppm (besides a short overshoot, see Fig. 2a), such that the ON state continues to exist. By contrast, the SSP2-4.5 and SSP4-6.0 scenarios stay above 460 ppm after the year 2050, such that we assume the model is in the monostable regime from then onwards.
@@ VI.2 Stochastic bifurcation in the GISS model
The diverging AMOC behaviour of ensemble members observed in PlaSim-LSG under SSP2-4.5 is intriguingly reminiscent of the so-called “stochastic bifurcation” found under the same forcing scenario in the CMIP6 model GISS-E2-1-G [52]. Can the global stability perspective presented here help explain the dynamics in that more comprehensive model?
@@
We select three ensemble members of the GISS model simulations under SSP2-4.5 (members r1i1p1f2, r7i1p1f2, and r10i1p1f2). The simulations extend until the year 2500 and show divergent AMOC behaviour (Fig. 14c). Following an initial AMOC weakening in all members, the first member starts to recover around the year 2100, the second recovers after 2200, whereas the third remains in a weak AMOC state until 2500 (but eventually recovers, see Ref. [52]). Similar to PlaSim-LSG, the weak AMOC state is characterised by a collapse of the overturning cell north of 45∘N while a weak overturning circulation is maintained south of 45∘N (Fig. 14h). In the year 1995 CE, when CO2 levels are at 360 ppm and the AMOC is in the ON state, the AMOC is around 30% stronger in GISS than in PlaSim-LSG, though the meridional streamfunctions have a qualitatively similar shape (Fig. 14e, g)).
@@
To relate these simulations to our results, we likewise select three ensemble members from the SSP2-4.5 simulations with PlaSim-LSG, based on their qualitative similarity with each of the GISS ensemble member (Fig. 14a). The first member maintains a strong AMOC, the second undergoes a weakening to about 5 Sv followed by a recovery, and the third collapses to the OFF state. Note that the evolution of the trajectories is delayed in PlaSim-LSG compared to the GISS model, and the initial AMOC weakening is less pronounced.
@@
We now compare the reduced state space trajectories of these simulations between the two models. Using the same definitions as for PlaSim-LSG, we compute the meridional and vertical SG in GISS based on the Atlantic zonally averaged salinity field. Due to the complexity of the GISS model, its AMOC stability landscape with respect to CO2 and the properties of potential edge states or ghost states are not known. However, we can study how the trajectories relate to the edge state and ghost state found in PlaSim-LSG.
@@
Remarkably, the reduced state space dynamics are qualitatively similar between the two models (Fig. 14b, d). The GISS trajectories start off from a significantly lower meridional SG than the PlaSim-LSG trajectories, in line with the fact that the AMOC in GISS is significantly stronger at that time. The vertical SG values are in good agreement between the models. As CO2 forcing increases, all trajectories initially move towards larger meridional and slightly larger vertical SG values. The AMOC recovery in GISS is characterised by a counter-clockwise rotation, where the loop performs a larger excursion for the trajectory that recovers from lower AMOC values. Interestingly, the reversal of the GISS trajectory with a late recovery occurs directly in the state space region where the edge state is located in PlaSim-LSG, following a path that resembles that of the recovering PlaSim-LSG trajectory. The collapsing trajectory in GISS skims the bottom end of the ghost state region before traveling to high meridional and low vertical SG values. This path is qualitatively similar to the collapsing PlaSim-LSG trajectory. Although the collapsing GISS simulation does not display any AMOC oscillations seen in the collapsing PlaSim-LSG simulation, there are still upward spikes in the state space trajectory that might hint at similar, yet dampened dynamics.
@@
To summarise, we find that the splitting of the GISS ensemble occurs in the same region of the projected state space in which the edge state is located in PlaSim-LSG. This supports the proposition that the “stochastic bifurcation” could indeed be a signature of a chaotic AMOC edge state near a boundary crisis.
@@ VII Discussion and conclusion
Traditionally, studies of climate tipping points often focus on the local dynamics near stable equilibria. Particularly, statistical EWS based on critical slowing-down measure changes in the local stability of an attractor under a quasi-adiabatic parameter drift: as the system approaches a bifurcation, the quasipotential flattens around the attractor, implying a reduction in restoring forces. Another way to look at this is that the barrier imposed by the edge state diminishes towards the bifurcation, and that the quasipotential flattens around the edge state (with an opposite sign of the curvature). This fits to our observation that edge tracking was more expensive at lower CO2 concentrations, further from the boundary crisis. Closer to the crisis, the edge state becomes “stickier” [105, 54] in the sense that trajectories tend to spend longer times in its vicinity, suggesting an alternative, non-local angle on critical slowing-down333Ying-Cheng Lai, personal communication.. It seems clear that the current rate of anthropogenic greenhouse gas emissions is forcing the climate system out of a steady-state. Our results indicate that the edge state dynamics can become relevant under plausible future emissions scenarios: the boundary crisis in PlaSim-LSG occurs at CO2 levels that could be reached within two decades. Nonetheless, whether this boundary crisis is a feature of the real climate system remains unknown.
@@
This study was conducted with a climate model of intermediate complexity that inevitably relies on simplifications and neglects numerous processes of potential relevance. Therefore, the results of our investigation may be highly model-dependent and not representative of reality. Even though the AMOC edge state found in PlaSim-LSG is a physically sensible steady-state, its nonlinear dynamics might be exaggerated effects of the highly simplified parameterisations of, e.g., sea ice and oceanic convection. On the other hand, we believe that PlaSim-LSG is to date the most complex climate model in which an edge state has been explicitly computed. Our results thus add a significant step towards realism to recent studies investigating edge states of the AMOC in a conceptual climate model [48] and a global ocean-only circulation model [57]. Furthermore, the similarity in the dynamics between PlaSim-LSG and the more complex GISS model suggests that the global stability view established here could provide key insights into the behaviour of state-of-the-art earth system models. As is increasingly clear, AMOC metastability and tipping behaviour is not restricted to simple climate models but occurs across the model hierarchy [52, 33, 79, 107]. The state space and parameter space of large models are just more challenging to explore.
@@
The key limitation for applying the edge tracking algorithm in even higher-dimensional systems is the computational cost of running long simulations. In PlaSim-LSG, producing one year of edge trajectory required on average 50 (90) years of simuation at (460 ppm). Thus, the 1400 year-long edge trajectory in the case consumed around 70 000 simulation years or 3000 CPU hours. Of course, this number depends strongly on the system and could be optimised by tuning the settings of the edge tracking procedure.
@@
The edge tracking algorithm converges to an edge state despite the complex geometry of the basin boundary, which is typically fractal [63, 10, 48]. In our study, we observed that the basin boundary is folded along transects of interpolated initial conditions, hinting at fractality. Due to the long lifetime of the ghost state, edge tracking appears to work for multiple iterations even beyond the boundary crisis (in the monostable regime). This permits probing ghost states while also demonstrating the difficulty of precisely determining critical forcing thresholds [48]. From a modelling perspective, it is not obvious that the interpolation between initial conditions in all dynamical variables yields new initial conditions that generate numerically stable and physical trajectories. We argue that convexity of the equations governing the climate dynamics ensures that trajectories quickly relax to a physical state.
@@
Using an ocean-only model, Ref. [57] found that the AMOC edge state features a less “spicy” (i.e., colder and fresher) deep North Atlantic than the attractors, as well as a higher dynamic enthalpy. Our results corroborate this in a coupled climate model while also revealing a much richer dynamics due to the ice-ocean-atmosphere coupling, in which the upper ocean plays a more active role. Whereas Ref. [57] concluded that the most relevant regions for anticipating AMOC transitions are located in the deep sea, our results suggest that many excursive observables are found also in the surface ocean. This could potentially be exploited for improved early warning systems of AMOC changes [100, 108].
@@
AMOC oscillations have received wide interest due to their occurrence in various climate models and potential for explaining past abrupt climate change. In the context of Dansgaard-Oeschger events [3], previous work has determined “sweet spots” for oscillations in parameter space [109]. We demonstrate a sweet spot in state space: while the ON and OFF states do not exhibit oscillations at , oscillations occur near the edge state. The drivers of these unstable oscillations involve similar processes previously identified in stable oscillation mechanisms in other models [110, 111, 112]. Dynamically, the presence of unstable oscillations near the crisis might hint at the existence of a subcritical Hopf bifurcation with respect to CO2. Oscillations in the AMOC strength and other popular observables such as the freshwater transport into the Atlantic further suggest that such quantities may be poor indicators of AMOC stability in out-of-equilibrium conditions.
@@
The limited predictability of the AMOC near an instability has already been suggested by Ref. [49]. Refs. [50, 52] have recently reiterated this idea by demonstrating an ensemble splitting caused by internal variability under identical time-dependent forcing. Our findings allow to understand this behaviour in terms of an edge state and, beyond the boundary crisis, a ghost state. We can thus directly link the dynamics of earth system models to fundamental concepts of dynamical systems theory that are often only explored in low-dimensional systems.
Data availability.
Selected simulation data and source code for implementing the edge tracking algorithm in PlaSim-LSG are available at https://doi.org/10.5281/zenodo.17053348. Further raw model simulation output can be provided by the authors upon request.